### v1

# Differential Communication in Channels with Mobility and Delay Spread using Zak-OTFS

Abstract
Zak-transform based orthogonal time frequency space (Zak-OTFS) is a delay-Doppler (DD) domain modulation scheme in which the signal processing is carried out in the DD domain. The channel when viewed in the DD domain is predictable. However, even with Zak-OTFS, pilots need to be sent periodically, albeit at a lower rate. In this paper, we propose a differential communication scheme for Zak-OTFS systems that alleviates the need for periodic pilot transmission. Towards this, we analytically show that the detected data can be used as a pilot and that the channel estimate obtained from the detected data can enable further detection enabling the “differential” aspect of the communication. Specifically, we leverage the prediction capability of the DD channel in Zak-OTFS to use the channel estimate (obtained from detected data symbols treated as pilots) in the previous instant to detect data in the next instant and propagate this forward. The advantages are two fold. First, it allows the data symbols to enjoy higher energy since the energy that would otherwise be required for pilot symbols can also be allocated to data symbols. Second, it allows for full spectral efficiency compared to point or embedded pilots. Comparison with the full spectral efficiency achieving spread pilot scheme shows that the proposed method achieves better bit-error rate at lower complexity.
Index Terms:
Channel predictability, delay-Doppler modulation, differential communication, pilot transmission, Zak-OTFS.
I Introduction
In wireless communication systems, information signals are communicated through a wireless medium. These signals undergo various distortions before being received at the receiver [1]. To estimate these distortions, the transmitter transmits reference signals called pilots, which are known to the receiver. Typically, these pilot symbols are transmitted at higher energy than the data symbols. The receiver carries out channel estimation, that is, an estimate of the channel is obtained using the received pilot symbols. For a rapidly varying channel, these pilot symbols are transmitted frequently. However, the pilot symbols do not carry any information and therefore frequent pilot transmissions lead to spectral efficiency loss. Different techniques are needed to reduce the frequency of pilot transmissions. Many schemes have been proposed to achieve this [2, 3, 4, 5]. In this paper we focus on the differential communication scheme.
Differential communication has been studied before in the context of space time block codes (STBC) [3, 4, 5]. In [3] the authors use an initial pilot transmission to estimate the channel. Next, assuming that the channel is constant for the next data transmission, the initial channel estimate is used to detect the STBC encoded data. The detected data is modeled as a pilot and used to estimate the channel again which enables more data detection. In [4], the authors use differential encoding of information symbols to enable detection at the receiver without the need for a channel estimate. Authors in [5] demonstrate that in addition to the advantages of pilot free communication, differential communication also provides rate advantages at higher layers above the physical layer. In the STBC context, differential communication is possible thanks to the channel becoming predictable in a multi-antenna system [5]. However, this is not readily applicable for single antenna system in a doubly selective channel. New modulation schemes are required for achieving predictability in such communication systems.
Recently, the authors in [6, 7] proposed a delay-Doppler (DD) domain modulation scheme called Zak-transform based orthogonal time frequency space (Zak-OTFS). One of the salient features of Zak-OTFS is that the channel when viewed in DD domain is predictable. This is because the channel changes as fast as the physics of the reflectors would allow, which in practice is slowly varying. With each Zak-OTFS frame typically spanning about few milliseconds, the channel remains almost stationary for a few frame transmission durations.
For channel estimation in Zak-OTFS many DD pilot frames have been proposed [7, 8, 9]. These include the point pilot (PP) frame [7], embedded pilot (EP) frame [8], and spread pilot (SP) frame [9]. The PP frame has a single non-zero value corresponding to the pilot in the whole frame. The EP frame has both pilot and data symbols separated by a guard region to prevent interference between the two. The SP frame has pilot symbols superimposed on the data symbols in way that makes the two mutually unbiased. The PP frame has the least spectral efficiency while SP has full spectral efficiency with EP frame in between. On the other hand, estimation complexity is the highest for SP frame followed by EP and PP frames.
In this paper, we propose differential communication scheme for doubly-dispersive channel using Zak-OTFS. Our contributions can be summarized as below.
  • We leverage the predictability of the DD domain channel in Zak-OTFS to design a communication scheme that reduces the frequency of pilot transmissions.
  • We propose a differential communication scheme in doubly-dispersive channels using Zak-OTFS. Since all practical channels are doubly-dispersive to various degrees, the method proposed in this paper has practical implications.
  • We analytically show that the detected data symbols can be used as pilot symbols to estimate the DD channel. We use a time-varying DD channel to show that the estimated channel in a given instant can enable data detection in the subsequent time instant and this process can be repeated over multiple time instants, alleviating the need for frequent pilot transmissions.
  • The proposed scheme makes for a low-complexity receiver while achieving full spectral efficiency and numerical results demonstrate that the proposed differential communication scheme achieves better bit-error performance when compared to full spectral efficiency achieving SP frame at lower complexity.
[…]
II Preliminaries
II-A Zak-OTFS
In Zak-OTFS, each information symbol is mounted on a pulsone which is a pulse train modulated by a tone [6, 7]. A Zak-OTFS frame consists of $MN$ DD bins, where $M$ is the number of delay bins and $N$ is the number of Doppler bins. For a data frame, $MN$ information symbols drawn from a constellation alphabet (e.g., 4-QAM) is mounted on these $MN$ DD bins by modulating each of the $MN$ pulsones. The input-output relation of the Zak-OTFS system [7] can be expressed as (we avoid providing a repeat of the system model for brevity) [^1] In this paper we consider the system model in discrete baseband and quantization and synchronization errors are not considered. However evaluating the performance of the system under these practical non-idealities is an important direction of future research.:
$\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n},$ (1)
where $\mathbf{y}\in\mathbb{C}^{MN\times 1}$ is the vector of received symbols in the DD domain, $\mathbf{H}\in\mathbb{C}^{MN\times MN}$ is the end-to-end channel matrix, $\mathbf{x}\in\mathbb{C}^{MN\times 1}$ is the vector of transmitted symbols in the time-domain, and $\mathbf{n}\in\mathbb{C}^{MN\times 1}$ is the additive Gaussian noise. For detection of $\mathbf{x}$ from $\mathbf{y}$, we need to compute an estimate of $\mathbf{H}$. This is carried out by transmitting pilots.
II-B Channel Estimation
To estimate the channel in a Zak-OTFS system, at the receiver, a DD domain cross-ambiguity function (which also happens to be the maximum-likelihood estimate of the channel [9]) is computed between $\mathbf{y}$ and $\mathbf{x}$. The DD domain cross-ambiguity function is given by:
$\mathbf{A}_{\mathbf{y},\mathbf{x}}[k,l]=\sum_{k^{\prime}=0}^{M-1}\sum_{l^{\prime}=0}^{N-1}\mathbf{Y}[k^{\prime},l^{\prime}]\mathbf{X}^{*}[k^{\prime}-k,l^{\prime}-l]e^{-\frac{j2\pi}{MN}l(k^{\prime}-k)},$ (2)
where $\mathbf{Y}[k,l]=\mathbf{y}[k+lM]$ and $\mathbf{X}[k,l]=\mathbf{x}[k+lM]$ are the matrix representations of the corresponding vectors, where $k=0,1,\cdots,M-1,l=0,1,\cdots,N-1$. The estimation of the channel matrix, $\mathbf{H}$ happens in two steps. First, $\mathbf{A}_{\mathbf{y},\mathbf{x}}$ gives the “model-free” [9] estimate, $\widehat{\mathbf{h}}_{\mathrm{eff}}$, of the effective channel matrix $\mathbf{h}_{\mathrm{eff}}$. Next, the estimate $\widehat{\mathbf{h}}_{\mathrm{eff}}\in\mathbb{C}^{M\times N}$ is used to construct the estimated channel matrix $\widehat{\mathbf{H}}$ (see [7, Eq. (38)]). $\widehat{\mathbf{H}}$ is then used for subsequent data detection.
II-C Ambiguity Function of PP
As mentioned earlier, in Zak-OTFS, information symbols are mounted on pulsone bases in the DD domain. The discrete DD domain point pulsone indexed by ($k_{0},l_{0}$) is:
$\mathbf{X}_{k_{0},l_{0}}^{(\mathrm{p})}[k,l]=\sum_{n,m\in\mathbb{Z}}e^{\frac{j2\pi}{N}nl}\delta[k-k_{0}-nM]\delta[l-l_{0}-mN],$ (3)
where the term $e^{\frac{j2\pi}{N}nl}$ is to account for quasi-periodicity of the pulsone [9]. Information symbols modulate pulsone at each $(k_{0},l_{0})$ tuple where $k_{0}=0,1,\cdots,M-1,l_{0}=0,1,\cdots,N-1$. The cross-ambiguity between a pulsone indexed by $(k_{0},l_{0})$ and $(k_{1},l_{1})$ is evaluated as (we skip writing the limits for the sum wherever it is clear, for brevity):
$\mathbf{A}_{\mathbf{X}_{k_{0},l_{0}}^{(\mathrm{p})},\mathbf{X}_{k_{1},l_{1}}^{(\mathrm{p})}}[k,l]$ $=\sum_{k_{1},l_{1}}\sum_{n_{1},m_{1}\in\mathbb{Z}}\sum_{n_{2},m_{2}\in\mathbb{Z}}e^{\frac{j2\pi}{N}n_{1}l^{\prime}}\times$
$\hskip 14.22636pt\delta[k^{\prime}-k_{0}-n_{1}M]\delta[l^{\prime}-l_{0}-m_{1}N]\times$
$\hskip 14.22636pte^{-\frac{j2\pi}{N}n_{2}(l^{\prime}-l)}\delta[k^{\prime}-k-k_{1}-n_{2}M]\times$
$\hskip 14.22636pt\delta[l^{\prime}-l-l_{1}-m_{2}N]e^{-\frac{j2\pi}{MN}l(k^{\prime}-k)}.$ (4)
From $\delta[k^{\prime}-k_{0}-n_{1}M]\delta[l^{\prime}-l_{0}-m_{1}N]$, we have $n_{1}=m_{1}=0$ and $k^{\prime}=k_{0},l^{\prime}=l_{0}$, since $k_{0},k^{\prime}=0,1,\cdots,M-1$ and $l_{0},l^{\prime}=0,1,\cdots,N-1$. Substituting in (4):
$\mathbf{A}_{\mathbf{X}_{k_{0},l_{0}}^{(\mathrm{p})},\mathbf{X}_{k_{1},l_{1}}^{(\mathrm{p})}}[k,l]$ $=\sum_{k_{1},l_{1}}\sum_{n_{2},m_{2}\in\mathbb{Z}}e^{-\frac{j2\pi}{N}n_{2}(l_{0}-l)}\times$
$\hskip 14.22636pt\delta[k_{0}-k-k_{1}-n_{2}M]\times$
$\hskip 14.22636pt\delta[l_{0}-l-l_{1}-m_{2}N]e^{-\frac{j2\pi}{MN}l(k_{0}-k)}.$ (5)
$\delta[k-(k_{0}-k_{1})-n_{2}M]\implies k=(k_{0}-k_{1})\bmod M$ and $\delta[l-(l_{0}-l_{1})-m_{2}N]\implies l=(l_{0}-l_{1})\bmod N$. This means that the cross-ambiguity is supported (alternately, the sum is non-zero) on a lattice given by the points $((k_{0}-k_{1})_{M},{(l_{0}-l_{1})_{N}})$. The self ambiguity $\mathbf{A}_{\mathbf{X}^{(p)},\mathbf{X}^{(p)}}[k,l]$ of the pulsone is supported on the lattice $(nM,mN)$ [9] and therefore (up to a phase):
$\mathbf{A}_{\mathbf{X}_{k_{0},l_{0}}^{(\mathrm{p})},\mathbf{X}_{k_{1},l_{1}}^{(\mathrm{p})}}[k,l]=\mathbf{A}_{\mathbf{X}^{(\mathrm{p})},\mathbf{X}^{(\mathrm{p})}}[k-(k_{0}-k_{1}),l-(l_{0}-l_{1})].$ (6)
II-D Channel Realization in DD Domain
To generate the time-varying DD domain channel, we make use of the following equations. For $i$th path ($i=0,1,\cdots,P-1$, for a $P$ path channel) with delay $\tau_{i}$ and Doppler spread $\nu_{i}$, the distance between the transmitter and receiver is:
$d_{i}(t)=\tau_{i}c+\frac{\nu_{i}c}{f_{c}}t,$ (7)
where $f_{c}$ is the carrier frequency and $c$ is the speed of light. To generate the channel gains we assume a path loss model given by:
$h_{i}(t)=\frac{\alpha_{i}}{d_{i}(t)}e^{j\theta},$ (8)
where $\alpha_{i}$ is a function of power profile of the channel and $\theta\sim\mathrm{U}[-2\pi,2\pi)$. For a given time $t$, the tuple $(h_{i}(t),\tau_{i},\nu_{i})$ characterizes the DD channel.
III Differential Communication in Zak-OTFS
In this section, we derive the conditions that enable differential communication in Zak-OTFS. The time-domain symbols mounted on pulsone bases is:
$\mathbf{x}[n]=\sum_{k_{0},l_{0}}\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}}[n]\mathbf{X}[k_{0},l_{0}],$ (9)
where $\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}}[n]$ is the $(k_{0},l_{0})$th pulsone basis in the time-domain (obtained by the inverse discrete Zak transform [10] of (3)), $\mathbf{X}[k_{0},l_{0}]$ is the $(k_{0},l_{0})$th data symbol mounted on the corresponding pulsone bases. The input output relation in time-domain is given by [11]:
$\mathbf{y}[n]=\sum_{k_{0},l_{0}}\sum_{k,l}$ $\mathbf{h}_{\mathrm{eff}}[k,l]\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}}[n-k]e^{\frac{j2\pi}{MN}l(n-k)}\mathbf{X}[k_{0},l_{0}]$
[…]
The maximum likelihood estimate of the channel is equivalent to computing the cross-ambiguity between the received and transmitted symbols [9, Appendix H]. Therefore, at the receiver, to estimate the channel, we compute the time-domain cross-ambiguity between the received time-domain symbols (in (10)) and the transmitted symbols (in (9)):
$\mathbf{A}_{\mathbf{y},\mathbf{x}}[k^{\prime},l^{\prime}]$ $=\frac{1}{MN}\sum_{n=0}^{MN-1}\mathbf{y}[n]\mathbf{x}^{*}[n-k^{\prime}]e^{-\frac{j2\pi}{MN}l^{\prime}(n-k^{\prime})}$
$=\frac{1}{MN}\Bigg{(}\sum_{n=0}^{MN-1}\sum_{k_{0},l_{0}}\sum_{k,l}\mathbf{h}_{\mathrm{eff}}[k,l]\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}}[n-k]\times$
$\hskip 14.22636pte^{\frac{j2\pi}{MN}l(n-k)}\mathbf{X}[k_{0},l_{0}]\sum_{k_{1},l_{1}}(\mathbf{x}^{(\mathrm{p})}_{k_{1},l_{1}}[n-k^{\prime}])^{*}\times$
$\hskip 14.22636pt\mathbf{X}^{*}[k_{1},l_{1}]e^{-\frac{j2\pi}{MN}l^{\prime}(n-k^{\prime})}+$
$\hskip 14.22636pt\sum_{n=0}^{MN-1}\mathbf{n}[n]\mathbf{x}^{*}[n-k^{\prime}]e^{-\frac{j2\pi}{MN}l^{\prime}(n-k^{\prime})}\Bigg{)}$
=$\overset{\small(a)}{=}\frac{1}{MN}\sum_{k,l}\mathbf{h}_{\mathrm{eff}}[k,l]\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\times$
$\hskip 14.22636pt\sum_{k_{1},l_{1}}\mathbf{X}^{*}[k_{1},l_{1}]\sum_{n=0}^{MN-1}\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}}[n-k]e^{\frac{j2\pi}{MN}l(n-k)}\times$
$\hskip 14.22636pt(\mathbf{x}^{(\mathrm{p})}_{k_{1},l_{1}}[n-k^{\prime}])^{*}e^{-\frac{j2\pi}{MN}l^{\prime}(n-k^{\prime})},$ (11)
where step ${\small(a)}$ follows because the pulsone samples and noise samples are uncorrelated. Substituting $\bar{n}=n-k$ in (11):
$\mathbf{A}_{\mathbf{y},\mathbf{x}}[k^{\prime},l^{\prime}]$ $=\frac{1}{MN}\sum_{k,l}\mathbf{h}_{\mathrm{eff}}[k,l]\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\times$
$\hskip 14.22636pt\sum_{k_{1},l_{1}}\mathbf{X}^{*}[k_{1},l_{1}]\sum_{\bar{n}=-k}^{MN-1-k}\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}}[\bar{n}]e^{\frac{j2\pi}{MN}l\bar{n}}\times$
$\hskip 14.22636pt(\mathbf{x}^{(\mathrm{p})}_{k_{1},l_{1}}[\bar{n}-(k^{\prime}-k)])^{*}e^{-\frac{j2\pi}{MN}l^{\prime}(\bar{n}-(k^{\prime}-k))}$
$=\sum_{k,l}\mathbf{h}_{\mathrm{eff}}[k,l]\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\sum_{k_{1},l_{1}}\mathbf{X}^{*}[k_{1},l_{1}]\times$
$\hskip 14.22636pt\mathbf{A}_{\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}},\mathbf{x}^{(\mathrm{p})}_{k_{1},l_{1}}}[k^{\prime}-k,l^{\prime}-l]e^{\frac{j2\pi}{MN}l^{\prime}(k^{\prime}-k)}$
$=\mathbf{h}_{\mathrm{eff}}[k^{\prime},l^{\prime}]\ \ *_{\sigma}$
$\underbrace{\left(\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\sum_{k_{1},l_{1}}\mathbf{X}^{*}[k_{1},l_{1}]\mathbf{A}_{\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}},\mathbf{x}^{(\mathrm{p})}_{k_{1},l_{1}}}[k^{\prime},l^{\prime}]\right)}_{\mathbf{B}[k^{\prime},l^{\prime}]}.$ (13)
Since $\mathbf{A}_{\mathbf{x}^{(\mathrm{p})}_{k_{0},l_{0}},\mathbf{x}^{(\mathrm{p})}_{k_{1},l_{1}}}[k^{\prime},l^{\prime}]=\mathbf{A}_{\mathbf{X}_{k_{0},l_{0}}^{(\mathrm{p})},\mathbf{X}_{k_{1},l_{1}}^{(\mathrm{p})}}[k,l]$ (i.e., the time-domain ambiguity is same as the DD domain ambiguity), substituting (6) in $\mathbf{B}[k^{\prime},l^{\prime}]$, we have:
$\mathbf{B}[k,l]$ $=\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\sum_{k_{1},l_{1}}\mathbf{X}^{*}[k_{1},l_{1}]\times$
$\hskip 14.22636pt\mathbf{A}_{\mathbf{X}^{(\mathrm{p})},\mathbf{X}^{(\mathrm{p})}}[k-(k_{0}-k_{1}),l-(l_{0}-l_{1})].$ (14)
Note $\mathbf{A}_{\mathbf{X}^{(\mathrm{p})},\mathbf{X}^{(\mathrm{p})}}[k-(k_{0}-k_{1}),l-(l_{0}-l_{1})]=\mathds{1}_{\{k=(k_{0}-k_{1})\bmod M,l=(l_{0}-l_{1})\bmod N\}}$. Substituting $k_{1}=(k-k_{0})_{M}$ and $l_{1}=(l-l_{0})_{N}$ in (14):
$\mathbf{B}[k,l]$ $=\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\mathbf{X}^{*}[(k-k_{0})_{M},(l-l_{0})_{N}].$ (15)
This implies that the cross-ambiguity between the transmitted data frame and the received frame is the twisted convolution between the effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ and the inner product between the transmitted data and its delay and Doppler shifted version. If the information symbols are chosen uniformly at random from a constellation, asymptotically we have:
$\mathbf{A}_{\mathbf{y},\mathbf{x}}[k^{\prime},l^{\prime}]\approx\mathbf{h}_{\mathrm{eff}}[k^{\prime},l^{\prime}]*_{\sigma}e_{\mathrm{d}}\delta[k]\delta[l]=e_{\mathrm{d}}\mathbf{h}_{\mathrm{eff}}[k^{\prime},l^{\prime}],$ (16)
where $e_{\mathrm{d}}$ is the energy of information symbols. The cross-ambiguity between the received and transmitted data symbols is therefore approximately equal to the estimate of the effective channel up to a scale.
Remark 1: Notice that in the case of spread pilot ([9] or [12]), $\mathbf{B}[k,l]$ is exactly $\delta[k]\delta[l]$ and therefore the cross-ambiguity $\mathbf{A}_{\mathbf{y},\mathbf{x}}[k,l]$ provides the exact estimate of $\mathbf{h}_{\mathrm{eff}}[k,l]$.
IV Numerical Results
(a) Actual $\mathbf{h}_{\mathrm{eff}}$
(b) $\mathbf{h}_{\mathrm{eff}}$ estimated using SP [12]
(c) $\mathbf{h}_{\mathrm{eff}}$ estimated using data
Figure 1: Demonstrating differential communication scheme using data symbols as pilots. Perfect channel and channel estimated using pilots are added for reference. Noiseless system. Zak-OTFS system with $M=31,N=37,\nu_{p}=30$ kHz, root raised cosine (RRC) pulse shaping with parameter $\beta_{\tau}=\beta_{\nu}=0.6$. For the DO frame, data symbols are drawn from a 4-QAM constellation.
(a) NMSE
(b) BER
Figure 2: Performance curves with proposed differential communication scheme, SP [9] with varying power distribution, separate sensing and communication, and perfect CSI for 4-QAM modulation. Zak-OTFS system with $M=31,N=37,\nu_{p}=30$ kHz, root raised cosine (RRC) pulse shaping with parameter $\beta_{\tau}=\beta_{\nu}=0.6$.
TABLE I: Table showing a comparison between various schemes in literature.
Method Complexity Operation at receiver Spectral efficiency
Differential communication (Ours) $\mathbf{\mathcal{O}(M^{2}N^{2})}$ Cross-ambiguity Full
Spread pilot [9][12] $\mathcal{O}(M^{2}N^{2})$ Cross-ambiguity Full
$\mathcal{O}(M^{2}N^{2})$ Pilot removal
Separate sensing and communication [7] $\mathcal{O}(M^{2}N^{2})$ Cross-ambiguity Half
We compare the complexity of the proposed differential communication approach against other approaches for Zak-OTFS in literature. The comparison is presented in Table I. The complexity associated with the proposed approach is $\mathcal{O}(M^{2}N^{2})$ which is incurred from the cross-ambiguity computation, while that for the spread pilot is twice this because of the extra step of pilot removal from the received frame. Using separate frame for sensing and communication incurs the same complexity as the proposed approach but has poor spectral efficiency since one frame is dedicated to the pilot transmission.
Figure 1 compares the estimate of effective channel matrix obtained from SP (Fig. 1b) and from data (Fig. 1c) using the proposed approach against the actual channel (Fig. 1a). A noiseless system is considered for demonstration purposes. The estimate obtained using data has undulations while the estimate from SP is free from the non-ideality. This corroborates the derivation from (15), where for a data frame, $\mathbf{B}[k,l]$ is not a scaled delta function but for a SP frame it is (see Remark 1).
We provide the bit-error rate (BER) and normalized mean square error (NMSE) performance for the proposed differential communication scheme. For the simulations, we consider the Vehicular-A (VehA) channel [13] and generate channel realizations per Sec. II-D. The NMSE is computed as:
$\text{NMSE}=\frac{\|\widehat{\mathbf{h}}_{\mathrm{eff}}-{\mathbf{h}}_{\mathrm{eff}}\|_{2}^{2}}{\|{\mathbf{h}}_{\mathrm{eff}}\|_{2}^{2}},$ (17)
where $\widehat{\mathbf{h}}_{\mathrm{eff}}$ is the estimated effective channel matrix.
Note on SNR computations: We consider an SP frame with pilot energy $e_{\mathrm{p,dB}}=e_{\mathrm{d,dB}}-5$ dB where $e_{\mathrm{d,dB}}$ is the data energy in dB scale. The noise variance $\sigma^{2}=1$ and in the linear scale, $e_{\mathrm{d,lin}}=10^{\mathrm{SNR_{d}}/10}$, $\mathrm{SNR_{d}}$ is the data SNR. For the data-only (DO) frame, we consider $e_{\mathrm{do,lin}}=e_{\mathrm{p,lin}}+e_{\mathrm{d,lin}}$ to be the energy in the data symbols. Both the SP and DO frames, therefore, carry equal energy, for fair comparison. For all the DO simulations presented below, we transmit a pilot frame every 30 frame transmissions to curb error propagation.
IV-A Error Propagation
Figure 3: Instantaneous NMSE as a function of number of frames. DO frame with information symbols drawn from 4-QAM. For SNR = 0 dB, the instantaneous NMSE reduces every time a pilot frame is transmitted. For SNR = 25 dB, the number of errors is small and there is no significant error propagation.
Fig. 3 plots the instantaneous NMSE of the channel estimate obtained using the DO frame (using 4-QAM symbols) at the receiver at low and high SNR values. From the plot, it is observed that at low SNR (SNR = 0 dB) error propagation results in instantaneous NMSE increasing with the number of frames. It is also observed that every 30 DO frames, there is a dip in the NMSE value, wherein the estimate from the pilot frame curbs error propagation. However, at high SNR (SNR = 25 dB), the number of errors is so low that there is no significant error propagation between two pilot transmissions.
(a) NMSE
(b) BER
Figure 4: Performance curves with proposed differential communication scheme, SP [9] with varying power distribution, separate sensing and communication, and perfect CSI for 16-QAM modulation. Zak-OTFS system with $M=31,N=37,\nu_{p}=30$ kHz, root raised cosine (RRC) pulse shaping with parameter $\beta_{\tau}=\beta_{\nu}=0.6$.
IV-B 4-QAM
Figure 2 shows the NMSE and BER performance of the Zak-OTFS system for various schemes using 4-QAM modulation. For the SP frame we consider various data and pilot distributions parameterized by $\alpha$. Keeping the total energy in the frame to be $e_{\mathrm{do,lin}}$, we consider $e_{\mathrm{d,lin}}=\alpha e_{\mathrm{do,lin}}$ and $e_{\mathrm{p,lin}}=(1-\alpha)e_{\mathrm{do,lin}}$ for various $\alpha$ values. The NMSE performance is seen to improve as $\alpha$ is decreased from $0.9$ to $0.1$. This is expected as the energy in the pilot increases leading to better estimates. SP frame with 5 turbo iterations [14] achieves the best NMSE as it involves cycling between channel estimation and data detection. The separate sensing and communication scheme involves transmitting separate pilot and data frames and the estimate from the pilot frame is used to detect data in the data frame. The NMSE of the proposed receiver with the DO frame floors at $0.2$. Moving to the BER performance, separate sensing and communication achieves performance closest to that with perfect channel state information (CSI), followed by the SP frame with turbo iterations. The performance of the proposed receiver using the DO frame is better than SP frames with various $\alpha$ values. It is interesting to note that the performance with DO frame is the lower bound for the performance with SP frame.
Complexity: Here we compare the complexity qualitatively for full spectral efficiency achieving methods only. Separate sensing and communication does not achieve full spectral efficiency. The complexity is the highest for SP frame with turbo iterations owing to repeated channel estimation and data detection. For the SP frame without turbo iterations, there is an additional step involved to remove the contribution of pilot symbols before data is detected. The DO frame incurs the least complexity because there is no additional interference cancellation step. The DO frame therefore achieves better performance than SP at lower complexity.
IV-C 16-QAM
Figure 4 shows the NMSE and BER performance of the Zak-OTFS system with 16-QAM modulation. NMSE and BER trends are similar to that observed in 4-QAM. The SP frames with decreasing $\alpha$ values achieve better NMSE, while the DO frame achieves the best BER performance closest to the performance with perfect CSI, while being the least complex.
V Conclusions
In this paper, we proposed a novel differential communication scheme for Zak-OTFS systems. The proposed scheme leveraged the predictability of the DD channel in Zak-OTFS to reduce the frequency of pilot symbol transmissions. The channel estimate obtained at a time instant was used for data detection in the next time instant. This allowed the information symbols transmitted in a DO frame to be used as pilots thereby achieving full spectral efficiency. We analytically showed the the cross-ambiguity between the received data frame and transmitted data frame provided a model-free estimate of the channel. Simulations results showed that the proposed detector using the DO frame achieved good NMSE and BER performance. SP and DO frames both achieve full spectral efficiency, but the proposed scheme with the DO frame achieved better BER performance at lower complexity than the SP frame.